the lines y= 3+2/3x , y= -1/3x, and y= -4/3x+3 intersect to form a triangle. find the vertices of the triangle.
Answer by lwsshak3(11628)

the lines y= 3+2/3x , y= -1/3x, and y= -4/3x+3 intersect to form a triangle. find the vertices of the triangle.
(1) y= 3+2/3x
(2) y= -1/3x=-x/3
(3) y= -4/3x+3=-4x/3+3
..
Intersection of (1) and (2)
(1) y= 3+2/3x=3+2x/3
(2) y= -1/3x=-x/3
..
3+2x/3=-x/3
lcd=3
9+2x=-x
3x=-9
x=-3
y=-x/3=1
point of intersection: (-3,1)
..
Intersection of (1) and (3)
(1) y= 3+2/3x=3+2x/3
(3) y= -4/3x+3=-4x/3+3
..
3+2x/3=-4x/3+3
lcd=3
9+2x=-4x+9
6x=0
x=0
y=3+2x/3=3
point of intersection: (0,3)
..
Intersection of (2) and (3)
(2) y= -1/3x=-x/3
(3) y= -4/3x+3=-4x/3+3
..
-x/3=-4x/3+3
lcd=3
-x=-4x+9
3x=9
x=3
y=-x/3=-1
point of intersection: (3,-1)
..
vertices of the triangle: (-3,1), (0,3), (3,-1)